OpenStudy (anonymous):

Please help? I'll answer yours in return!!!! If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation.x + 6y + z = 2x - 3y + 2z = -1x + 5y + 3z = -4 Answer options: 11y + 4z = 62y + 5z = 3 y - 2z = 6-8y - z = 3 3y + 3z = -311y + 4z = -6 -9y + z = -3-y + 2z = -6

OpenStudy (anonymous):

(I think it's A...)

OpenStudy (anonymous):

x + 6y + z = 2 x - 3y + 2z = -1 x + 5y + 3z = -4 i guess those are the equations so isolate x from the first: x = 2-6y-z now plug x into the second and third : 2 - 6y -z - 3y +2z = -1 z - 9y = -3 2 - 6y -z + 5y + 3z = -4 2z -y = -6 so the two equations we have are : z - 9y = -3 2z -y = -6

OpenStudy (anonymous):

this is not A

OpenStudy (anonymous):

:(((((

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

i gave you the answer already .. now you should see which one is the same

OpenStudy (anonymous):

C or D? I don't see this as clearly as I should

OpenStudy (anonymous):

D -9y + z = -3 -y + 2z = -6 same as mine :)

OpenStudy (anonymous):

awwwww thank you kind sir